angle between the pair of straight lines 4x2 + 10xy + my2 + 5x + 10y = 0 is

Grade:12

2 Answers

Satyadip Mahapatra
8 years ago
general equation of pair of staight line –  ax+ 2hxy + by2 + 2gx + 2hy + c = 0 
so, a = 4, b = m, h = 5, g = 5/2, f = 5, c = 0.
put it in the equation – abc + 2hgf – af2 – bg2 – ch2 = 0
and get m = 4 
to get angle, put require value in the equation  
angle = tan -1 [ | 2 ( h2 – ab )1/2 | / | a + b | ]  
the answer will be  tan-1 ( ¾ ).
Satyadip Mahapatra
8 years ago
 
general equation of pair of staight line –  ax+ 2hxy + by2 + 2gx + 2hy + c = 0 
so, a = 4, b = m, h = 5, g = 5/2, f = 5, c = 0.
put it in the equation – abc + 2hgf – af2 – bg2 – ch2 = 0
and get m = 4 
to get angle, put require value in the equation  
angle = tan -1 [ | 2 ( h2 – ab )1/2 | / | a + b | ]  
the answer will be  tan-1 ( ¾ ).